OpenStudy (anonymous):

Using complete sentences, explain how you would factor completely \(x^9-27\)

OpenStudy (anonymous):

@Chlorophyll @jim_thompson5910 @phi @Hero

OpenStudy (anonymous):

@karatechopper @mattfeury @satellite73 @shadowfiend

OpenStudy (anonymous):

@tkhunny

OpenStudy (anonymous):

Formula a³ - b³

OpenStudy (anonymous):

keeping in mind that this has to be algebra 2 level...just sayin

OpenStudy (tkhunny):

For nearly everything that can be expressed like this, \[x^{n} - m^{n}\], (Certainly assuming x is Real and m and n are integers) it is true that (x-m) divides it. That was quite a sentence.

OpenStudy (anonymous):

\[x^9-27\]\[x^3\times x^3\times x^3-3^3\]\[\large{x^3}^3 -3^3\]\[\large\text{now what??}\]

OpenStudy (anonymous):

This proves that you have no clue about Formula a³ - b³ :P

OpenStudy (anonymous):

yes exactly :P ...im soooo confused :(

OpenStudy (anonymous):

a³ - b³ = ( a -b) ( a² + ab + b²) Do yourself a humongous favor, memorize it, will you!

OpenStudy (anonymous):

okay but i need to explain in sentences how to do this so....

OpenStudy (anonymous):

??

OpenStudy (anonymous):

1. Convert the terms into power of 3: x^9 = (x³)³ , 27 = 3³ 2. Plug into the formula ( I assume you know it by now ;) )

OpenStudy (anonymous):

wait wouldnt it just be \((x^3-3)^3\) ? \[\left((x^3-3)(x^3-3)(x^3-3)\right)\]

OpenStudy (anonymous):

(x³)³ - 3³ = ( x³ - 3 ) [ (x³)² + 3x³ + 3² ] = ....

OpenStudy (anonymous):

I LOVE YOU O_O

OpenStudy (anonymous):

urm....imma girl too.....so that was just sisterly love..LOL

OpenStudy (anonymous):

i got the help i needed, jeez :P ....what an oppressive being :P

OpenStudy (anonymous):

=)